OpenStudy (anonymous):

\[w = f(x,y)\\x(r, \theta) = r \cos \theta\\y(r, \theta) = r \sin \theta\] prove that \[{\partial{w} \over \partial{x}} = {\partial{w} \over \partial{r}} \cos \theta - {\partial{w} \over \partial\theta}{\sin \theta \over r}\]

OpenStudy (zarkon):

this is just an application of the chain rule...where are you stuck?

OpenStudy (zarkon):

use \[ {\partial{w} \over \partial{r}}= {\partial{w} \over \partial{x}} {\partial{x} \over \partial{r}}+ {\partial{w} \over \partial{y}} {\partial{y} \over \partial{r}}\] \[ {\partial{w} \over \partial{\theta}}= {\partial{w} \over \partial{x}} {\partial{x} \over \partial{\theta}}+ {\partial{w} \over \partial{y}} {\partial{y} \over \partial{\theta}}\] the follows quickly after that

OpenStudy (anonymous):

\[{\partial{w} \over \partial{x}}\] is where I am confused. I can't figure out how to take the partial of \[w = f(x, y)\]

OpenStudy (zarkon):

\[{\partial{w} \over \partial{r}} \cos \theta - {\partial{w} \over \partial\theta}{\sin \theta \over r}\] \[=\left[{\partial{w} \over \partial{x}} {\partial{x} \over \partial{r}}+ {\partial{w} \over \partial{y}} {\partial{y} \over \partial{r}}\right] \cos \theta - \left[{\partial{w} \over \partial{x}} {\partial{x} \over \partial{\theta}}+ {\partial{w} \over \partial{y}} {\partial{y} \over \partial{\theta}}\right]{\sin \theta \over r}\] \[=\cdots\]

OpenStudy (zarkon):

I assume you can find \[\frac{\partial x}{\partial r},\frac{\partial y}{\partial r},\frac{\partial x}{\partial \theta},\frac{\partial y}{\partial \theta}\]

OpenStudy (anonymous):

yes, the ones are easy for me

OpenStudy (anonymous):

those ones*

OpenStudy (zarkon):

ok...replace those in the expression I have above and simplify.

OpenStudy (anonymous):

ok, I'm working it out

OpenStudy (anonymous):

Okay, it all simplified down to: \[{\partial{w} \over \partial{x}} = {\partial{w} \over \partial{x}}\] It makes sense, just plug in the partials that I know how to find and it worked out!

OpenStudy (zarkon):

this is what I did... \[=\left[{\partial{w} \over \partial{x}} {\partial{x} \over \partial{r}}+ {\partial{w} \over \partial{y}} {\partial{y} \over \partial{r}}\right] \cos (\theta) - \left[{\partial{w} \over \partial{x}} {\partial{x} \over \partial{\theta}}+ {\partial{w} \over \partial{y}} {\partial{y} \over \partial{\theta}}\right]{\sin(\theta) \over r}\] \[=\left[{\partial{w} \over \partial{x}} \cos(\theta)+ {\partial{w} \over \partial{y}} \sin(\theta)\right] \cos (\theta) - \left[{\partial{w} \over \partial{x}} (-r\sin(\theta))+ {\partial{w} \over \partial{y}} r\cos(\theta)\right]{\sin (\theta) \over r}\] \[={\partial{w} \over \partial{x}} \cos^2(\theta)+ {\partial{w} \over \partial{y}} \sin(\theta) \cos \theta+ {\partial{w} \over \partial{x}} \sin^2(\theta))- {\partial{w} \over \partial{y}} \cos(\theta)\sin(\theta)\] \[={\partial{w} \over \partial{x}} \cos^2(\theta)+ {\partial{w} \over \partial{x}} \sin^2(\theta)={\partial{w} \over \partial{x}}\]

OpenStudy (anonymous):

Thanks for the help. That was only part a of the problem, I think that I can do the rest now

OpenStudy (zarkon):

yw